OpenStudy (anonymous):

d/dx lnsinx-(1/2)(sin^2)(x)

OpenStudy (anonymous):

\[\frac{d}{dx}\ln(\sin(x))-\frac{1}{2}\sin^2(x)\]??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok lots of chain rule here

OpenStudy (anonymous):

\[\frac{d}{dx}[\ln(x)]=\frac{1}{x}\] so \[\frac{d}{dx}[\ln(f(x))]=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

I know d/dx ln(sin(x)) is simplified to cos(x)/sin(x). Just not sure where to start on the second half

OpenStudy (anonymous):

I figured it out! Thank you!